OpenStudy (anonymous):

Let f(x)=(x−5)2. Find the average rate of change of f(x) with respect to x from x=a to x=a+h if h≠0. Simplify your answer as much as possible.

OpenStudy (anonymous):

Oops supposed to be (x-5)^2

myininaya (myininaya):

\[\frac{f(x_1)-f(x_2)}{x_1-x_2}\] You are given \[x_1=a+h, x_2=a\] Plug these in and simplify

myininaya (myininaya):

\[\frac{f(a+h)-f(a)}{a+h-a}\]

myininaya (myininaya):

what is f(a+h)?

OpenStudy (anonymous):

I have no idea sadly...

myininaya (myininaya):

Replace the x's in you function with a+h

myininaya (myininaya):